OpenStudy (anonymous):

.3+1(3)/(4) simplify

OpenStudy (anonymous):

Is it 0.3 or just 3?

OpenStudy (anonymous):

0.3

OpenStudy (anonymous):

its 4.2/4= 42/40

OpenStudy (anonymous):

0.3+3/4 (1.2+3)/4 =4.2/4

OpenStudy (anonymous):

It's 3.3 on the top. So 3.3/4 which is 0.825

OpenStudy (anonymous):

how is it 3.3 on the top , its 4.2 on the top

OpenStudy (anonymous):

thanks everyone, i am taking online algebra 1 to make up a credit to graduate so i am going to need alot of help :\

OpenStudy (anonymous):

Because 0.3 + 3 = 3.3

OpenStudy (anonymous):

no , the question is not (0.3 +3)/4 But 0.3 +(3/4)

OpenStudy (anonymous):

Well, the way it was written would imply the other way, but if you're right, then your answer is dead on.

OpenStudy (anonymous):

lol i am almost 100% sure its meant that way ,

OpenStudy (anonymous):

but am not gonna start messin with a Linux Wiz :D